OpenStudy (anonymous):

A recipe for molding dough calls for flour and salt in a ratio of 2 to 1. If you use 4 cups of salt, how much flour should you use? 2 4 5 8

OpenStudy (ohmaigosshh):

you would use 2 things of salt....

OpenStudy (ohmaigosshh):

you would use 2 cups of salt because flour is equal to 2 and salt is equal to 1..... so if u double the flour... making it 4, and double the salt, it would make the salt 2

OpenStudy (nurali):

2/1=x/4 8=x x=8